# Bose-Hubbard model with power-law hopping in one dimension

### v1

@@ Abstract
We characterize the SF-MI quantum phase transition at constant density $\rho=1$ by computing the mean-squared winding number fluctuations $\langle\mathcal{W}^{2}\rangle$ – a scale-invariant quantity proportional to the superfluid stiffness $Y_{s}$ as $\langle\mathcal{W}^{2}\rangle=Y_{s}/(LT)$ – up to sizes $L=512$. $\langle\mathcal{W}^{2}\rangle$ is expected to have a finite value and a zero value in the SF and MI phases, respectively. Figures 3(a) and (b) present example results for $\langle\mathcal{W}^{2}\rangle$ as a function of $t/U$ for two power-law exponents $\alpha<2$ and $\alpha>2$ in Eq. (1), respectively, and for different system sizes $L$. In the figures, the values of $\langle\mathcal{W}^{2}\rangle_{L_{\text{max}}}$ for the largest sizes $L_{\text{max}}$ used in the computations have been subtracted for clarity. Panels (a) and (b) show a clear crossing of $\langle\mathcal{W}^{2}\rangle$ vs $t/U$ when plotted for different $L$ at values $(t/U)_{c}=0.043\pm 0.0005$ and $(t/U)_{c}=0.131\pm 0.0008$, respectively. These crossings correspond to a quantum phase transition at the respective values of the critical ratio $(t/U)_{c}$ (see Insets for further extrapolation to thermodynamic limit). Interestingly, the very presence of a crossing in the $\langle\mathcal{W}^{2}\rangle-(t/U)$ curves rules out the Berezinskii-Kosterlitz-Thouless universality class for these power law models, in contrast to familiar short-range models in one dimension [10] and long-range models with power-law density-density interactions [39, 40, 10]. Similar crossings are observed for all $1<\alpha\leq 3$, implying a continuous scale-invariant phase transition in this whole range of $\alpha$. For $\alpha>3$ the transition is instead of the BKT type, consistent with short-range models. This fixes $\alpha_{*}=3$ for the model (1).
@@
Our predictions for the superfluid dispersion relation and the correlation functions can be directly measured in experiments with dipolar atoms and molecules ($\alpha=3$, [15, 47, 48, 49, 50]) and cold ions ($1<\alpha\lesssim 3$, [16, 22, 23]). Our work provides exact results to benchmark theories for long-range quantum models and opens up multiple other research directions, including the nature of the groundstate in higher dimensions.

### v2

@@ Abstract
We characterize the SF-MI quantum phase transition at constant density $\rho=1$ by computing the mean-squared winding number fluctuations $\langle\mathcal{W}^{2}\rangle$ – a scale-invariant quantity proportional to the superfluid stiffness $Y_{s}$ as $\langle\mathcal{W}^{2}\rangle=Y_{s}/(LT)$ – up to sizes $L=512$. $\langle\mathcal{W}^{2}\rangle$ is expected to have a finite value and a zero value in the SF and MI phases, respectively. Figures 3(a) and (b) present example results for $\langle\mathcal{W}^{2}\rangle$ as a function of $t/U$ for two power-law exponents $\alpha<2$ and $\alpha>2$ in Eq. (1), respectively, and for different system sizes $L$. In the figures, the values of $\langle\mathcal{W}^{2}\rangle_{L_{\text{max}}}$ for the largest sizes $L_{\text{max}}$ used in the computations have been subtracted for clarity. Panels (a) and (b) show a clear crossing of $\langle\mathcal{W}^{2}\rangle$ vs $t/U$ when plotted for different $L$ at values $(t/U)_{c}=0.0430\pm 0.0005$ and $(t/U)_{c}=0.131\pm 0.001$, respectively. These crossings correspond to a quantum phase transition at the respective values of the critical ratio $(t/U)_{c}$ (see Insets for further extrapolation to thermodynamic limit). Interestingly, the very presence of a crossing in the $\langle\mathcal{W}^{2}\rangle-(t/U)$ curves rules out the Berezinskii-Kosterlitz-Thouless universality class for these power law models, in contrast to familiar short-range models in one dimension [10] and long-range models with power-law density-density interactions [39, 40, 10]. Similar crossings are observed for all $1<\alpha\leq 3$, implying a continuous scale-invariant phase transition in this whole range of $\alpha$. For $\alpha>3$ the transition is instead of the BKT type, consistent with short-range models. This fixes $\alpha_{*}=3$ for the model (1).
@@
Our predictions for the superfluid dispersion relation and the correlation functions can be directly measured in experiments with dipolar atoms and molecules ($\alpha=3$, [15, 47, 48, 49, 50]) as well as cold ions ($1<\alpha\lesssim 3$, [16, 22, 23]) and are directly relevant to recent experiments with subradiant dipolar excitons in double quantum well systems [51]. Our work provides exact results to benchmark theories for long-range quantum models and opens up multiple other research directions, including the nature of the groundstate in higher dimensions.